OpenStudy (anonymous):

Which ordered pair is not a solution of y=-4x-9 A.(-9,-27) B.(3,-21) C.(-6,15) D.(5,-28)

OpenStudy (anonymous):

Someone please help I will give you a medal

OpenStudy (janu16):

y=-4+9 (x,y)=(-9,27) 27=-4*(-9)-9 27=36-9 27=27 so true

OpenStudy (janu16):

can you try other ones but pluggin those numbers? @nalah-445

OpenStudy (janu16):

no?

OpenStudy (faiqraees):

Still want help?

OpenStudy (janu16):

probably not

OpenStudy (anonymous):

Refer to the Mathematica attachment.